OpenStudy (anonymous):

(x+1/x)3 solve by indefinite integration

OpenStudy (amistre64):

expand it out and work the pieces

OpenStudy (anonymous):

by using (a+b)3 formula

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

\[\int\limits(x+\frac{ 1 }{ x })]dx =\[intxdx+\int1/x)dx =\[x^{2}/2+logx+c\]

OpenStudy (amistre64):

the posted problem has a ^3 attached to it. so you can just plow the integral thru like that.

OpenStudy (amistre64):

but once it is expanded out, that would be the process yes

OpenStudy (anonymous):

(x+1/x)^3=x^3+1/x^3+3(x+1/x) .X3+1/x^(3 ) +3*x*1/x(x+1/x)=x3+1/x^3 +3x+3/x=x3+x-3+3x+3*1/x ∫▒〖(x+1/x〗) 3dx=∫▒x^3 dx+∫▒x^(-3) dx+3∫▒x dx+3∫▒1/x dx = x^4/4+x^(-2)/(-2)+〖3x〗^2/2+3logx+c